# Partial Resilient Leader-Follower Consensus in Time-Varying Graphs

### v1

@@ IV-A Analysis of the BP-MSR Algorithm
We first define a set of all followers who are adjacent to at least $F+1$ normal leaders in the sequence $\mathbb{G}$ at any time $t\geq t^{\prime}$: $\mathcal{F}_{\mathcal{L}}=\{i_{1}\in\mathcal{F}_{\mathcal{C}}\mid\exists t_{1}\geq t^{\prime}\ \text{ s.t. }i_{1}\in\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{1}],\ |\mathcal{N}_{i_{1}}[t^{\prime}+t_{1}]\cap\mathcal{L}_{\mathcal{M}}|\geq F+1\}$. By definitions of strong $(2F+1)$-robustness and $F$-local model, for all $t\geq t_{C}$ such that $|\mathcal{F}_{\mathcal{R}}[t]|>0$, there must exist $i_{1}\in\mathcal{F}_{\mathcal{R}}[t]\cap\mathcal{F}_{\mathcal{L}}$. Since $\mathcal{F}_{\mathcal{C}}$ is finite and $\mathcal{F}_{\mathcal{L}}\subseteq\mathcal{F}_{\mathcal{C}}$, there must exist a finite $T_{1}\geq 1$ such that $\mathcal{F}_{\mathcal{L}}\subseteq\bigcup_{t_{1}=0}^{T_{1}-1}\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{1}]$. Therefore, for any $t_{1}\in[0,T_{1}-1]$ where $|\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{1}]|>0$, $\mathcal{S}_{1,t_{1}}=\{i_{1}\in\mathcal{F}_{\mathcal{C}}\mid\mathcal{N}_{i_{1}}[t^{\prime}+t_{1}]\cap\mathcal{L}_{\mathcal{M}}|\geq F+1\}$ is nonempty. Then, through the BP-MSR algorithm, follower $i_{1}$ will use at least one $C_{r}$ to update $x_{i_{1}}[t^{\prime}+t_{1}+1]$.
@@
Furthermore, by definition, $\mathcal{G}_{\mathcal{R}}[t]$ is strongly $(2F+1)$-robust with respect to $\mathcal{L}\cup\mathcal{A}$ by Lemma 2 at $t$ if $|\mathcal{F}_{\mathcal{R}}[t]|>0$. Therefore, for any $t_{2}\in[T_{1},T_{2}-1]$ where $|\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{2}]|>0$, $\mathcal{S}_{2,t_{2}}=\{i_{2}\in\mathcal{S}_{X}(t^{\prime}+t_{2},\alpha^{t_{2}}\underline{\epsilon},\alpha^{t_{2}}\overline{\epsilon})\mid|\mathcal{D}_{i_{2}}[t^{\prime}+t_{2}]|\geq 2F+1\}$ is nonempty if $\mathcal{S}_{X}(t^{\prime}+t_{2},\alpha^{t_{2}}\underline{\epsilon},\alpha^{t_{2}}\overline{\epsilon})\neq\emptyset$. If it is nonempty, since $\mathcal{A}$ is $F$-local, $\mathcal{D}_{i_{2}}[t^{\prime}+t_{2}]$ contains at least $F+1$ normal agents $\forall i_{2}\in\mathcal{S}_{2,t_{2}}$. Then, by definition of $\mathcal{S}_{X}(t^{\prime}+t_{2},\alpha^{t_{2}}\underline{\epsilon},\alpha^{t_{2}}\overline{\epsilon})$, $x_{i_{2}}[t^{\prime}+t_{2}]>x_{j}^{i_{2}}[t^{\prime}+t_{2}]$ or $x_{i_{2}}[t^{\prime}+t_{2}]<x_{j}^{i_{2}}[t^{\prime}+t_{2}]$ $\forall j\in\mathcal{D}_{i_{2}}[t^{\prime}+t_{2}]$. This means $i_{2}$ will use at least one in-neighbor’s state within $\overline{m}[t^{\prime}]+\alpha^{t_{2}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{t_{2}}\overline{\epsilon}]$ to update $x_{i_{2}}[t^{\prime}+t_{2}+1]$, whose bounds are $x_{i_{2}}[t^{\prime}+t_{2}+1]\in[\overline{m}[t^{\prime}]+\alpha^{t_{2}+1}\underline{\epsilon},\overline{M}[t^{\prime}]+\alpha^{t_{2}+1}\underline{\epsilon}]$ by (8) and (9), which is also contained by $\overline{m}[t^{\prime}]+\alpha^{T_{2}}\underline{\epsilon},\overline{M}[t^{\prime}]+\alpha^{T_{2}}\underline{\epsilon}]$. Now denote $\mathcal{S}_{2}=\cup_{t_{2}=T_{1}}^{T_{2}-1}\mathcal{S}_{2,t_{2}}$. Because (i) $\mathcal{S}_{2}\subseteq\mathcal{K}_{2}$, (ii) $\mathcal{K}_{1}\subseteq\mathcal{K}_{2}$, (iii) $\mathcal{S}_{2,T_{1}}\cap\mathcal{K}_{1}=\emptyset$, and (iv) $i_{2}\notin\mathcal{S}_{X}(t^{\prime}+T_{2},\alpha^{T_{2}}\underline{\epsilon},\alpha^{T_{2}}\overline{\epsilon})\ \forall i_{2}\in\mathcal{K}_{2}$, we get $|\mathcal{S}_{X}(t^{\prime}+T_{2},\alpha^{T_{2}}\underline{\epsilon},\alpha^{T_{2}}\overline{\epsilon})|<|\mathcal{S}_{X}(t^{\prime}+T_{1},\alpha^{T_{1}}\underline{\epsilon},\alpha^{T_{1}}\overline{\epsilon})|$.
@@
This logic of the existence of finite time $T_{p}\geq T_{p-1}+1$ such that $|\mathcal{S}_{X}(t^{\prime}+T_{p},\alpha^{T_{p}}\underline{\epsilon},\alpha^{T_{p}}\overline{\epsilon})|<|\mathcal{S}_{X}(t^{\prime}+T_{p-1},\alpha^{T_{p-1}}\underline{\epsilon},\alpha^{T_{p-1}}\overline{\epsilon})|$ can be continued iteratively for $p\geq 2$. We first define $\mathcal{K}_{p}=\{i_{p}\in\mathcal{F}_{\mathcal{R}}[t^{\prime}+T_{p}]\mid x_{i_{p}}[t^{\prime}+T_{p}]\in[\overline{m}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{T_{p}}\overline{\epsilon}]\}$. Furthermore, we define $\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]=\mathcal{N}_{i_{p}}[t^{\prime}+t_{p}]\cap\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{p}]\setminus\mathcal{S}_{X}(t^{\prime}+t_{p},\alpha^{t_{p}}\underline{\epsilon},\alpha^{t_{p}}\overline{\epsilon})$ for any $t_{p}\in[T_{p-1},T_{p}-1]$. Let $T_{p}\geq T_{p-1}+1$ be the earliest time where $\mathcal{K}_{p-1}\subseteq\mathcal{K}_{p}$. Using the prior arguments, we conclude that $x_{i_{p-1}}[t^{\prime}+T_{p}]\in[\overline{m}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon}]$ $\forall i_{p-1}\in\mathcal{K}_{p-1}$, and thus $\mathcal{K}_{p-1}\subseteq\mathcal{K}_{p}$. Because all $i_{p-1}\in\mathcal{K}_{p-1}\subseteq\mathcal{F}_{\mathcal{C}}$ must belong to $\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{p}]$ at some $t_{p}\in[T_{p-1},T_{p}-1]$, such finite $T_{p}$ exists.
@@
Furthermore, using the prior arguments, for any $t_{p}\in[T_{p-1},T_{p}-1]$ where $|\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{p}]|>0$, we can have nonempty $\mathcal{S}_{p,t_{p}}=\{i_{p}\in\mathcal{S}_{X}(t^{\prime}+t_{p},\alpha^{t_{p}}\underline{\epsilon},\alpha^{t_{p}}\overline{\epsilon})\mid|\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]|\geq 2F+1\}$. If it is nonempty, since $\mathcal{A}$ is $F$-local, $\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]$ contains at least $F+1$ normal agents $\forall i_{p}\in\mathcal{S}_{p,t_{p}}$. Then, by definition of $\mathcal{S}_{X}(t^{\prime}+t_{p},\alpha^{t_{p}}\underline{\epsilon},\alpha^{t_{p}}\overline{\epsilon})$, $x_{i_{p}}[t^{\prime}+t_{p}]>x_{j}^{i_{p}}[t^{\prime}+t_{p}]$ or $x_{i_{p}}[t^{\prime}+t_{p}]<x_{j}^{i_{p}}[t^{\prime}+t_{p}]$ $\forall j\in\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]$. This means $i_{p}$ will use at least one in-neighbor’s state within $\overline{m}[t^{\prime}]+\alpha^{t_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{t_{p}}\overline{\epsilon}]$ to update $x_{i_{p}}[t^{\prime}+t_{p}+1]$, whose bounds are $x_{i_{p}}[t^{\prime}+t_{p}+1]\in[\overline{m}[t^{\prime}]+\alpha^{t_{p}+1}\underline{\epsilon},\overline{M}[t^{\prime}]+\alpha^{t_{p}+1}\underline{\epsilon}]$ with the similar reasoning as (8) and (9), which is also contained by $\overline{m}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon}]$. Now denote $\mathcal{S}_{p}=\cup_{t_{p}=T_{p-1}}^{T_{p}-1}\mathcal{S}_{p,t_{p}}$. Then, because (i) $\mathcal{S}_{p}\subseteq\mathcal{K}_{p}$, (ii) $\mathcal{K}_{p-1}\subseteq\mathcal{K}_{p}$, (iii) $\mathcal{S}_{p,T_{p-1}}\cap\mathcal{K}_{p-1}=\emptyset$, and (iv) $i_{p}\notin\mathcal{S}_{X}(t^{\prime}+T_{p},\alpha^{T_{p}}\underline{\epsilon},\alpha^{T_{p}}\overline{\epsilon})\ \forall i_{p}\in\mathcal{K}_{p}$, we get $|\mathcal{S}_{X}(t^{\prime}+T_{p},\alpha^{T_{p}}\underline{\epsilon},\alpha^{T_{p}}\overline{\epsilon})|<|\mathcal{S}_{X}(t^{\prime}+T_{p-1},\alpha^{T_{p-1}}\underline{\epsilon},\alpha^{T_{p-1}}\overline{\epsilon})|$.
@@
$m[t^{\prime}]+\alpha^{\overline{T}}\underline{\epsilon}\leq x_{i}[t^{\prime}+\overline{T}]\leq M[t^{\prime}]-\alpha^{\overline{T}}\overline{\epsilon}.$
Considering V[t]=M[t]m[t]V[t]=M[t]-m[t], we get:
$V[t^{\prime}+\overline{T}]$ $=M[t^{\prime}+\overline{T}]-m[t^{\prime}+\overline{T}]$
$\leq M[t^{\prime}]-\alpha^{\overline{T}}\overline{\epsilon}-\left(m[t^{\prime}]+\alpha^{\overline{T}}\underline{\epsilon}\right)$
@@
Using the fact that $\underline{\epsilon}=C_{r}-m[t^{\prime}]$ and $\overline{\epsilon}=M[t^{\prime}]-C_{r}$, from (10), we get
@@
The proof of Theorem 1 follows the similar approach of [4, Theorem 1], but with a key distinction. While [4, Theorem 1] assumes that the entire graph is strongly $(2F+1)$-robust with respect to $\mathcal{L}$ over a fixed time period, our analysis assumes that some individual followers become active and thus belong to a strongly $(2F+1)$-robust subgraph infinitely often.
@@
The corollary shows that if there exists an infinite sequence of finite time instants $\{\tau_{i}\}_{i\in\mathbb{Z}_{\geq 0}}$ such that $\bigcup_{t=\tau_{i}}^{\tau_{i+1}}\mathcal{G}[t]$ has all normal followers activated via (1) (which implies that the union of graph is strongly $(2F+1)$-robust), then the normal followers can achieve full resilient leader-follower consensus. This result in a way generalizes [4, Theorem 1], which requires the entire network to be strongly $(2F+1)$-robust over every fixed window of $T$ time steps.

### v2

@@ IV-A Analysis of the BP-MSR Algorithm
We first define a set of all followers who are adjacent to at least $F+1$ normal leaders in the sequence $\mathbb{G}$ at any time $t\geq t^{\prime}$: $\mathcal{F}_{\mathcal{L}}=\{i_{1}\in\mathcal{F}_{\mathcal{C}}\mid\exists t_{1}\in\mathbb{Z}_{\geq 0}\ \text{ s.t. }i_{1}\in\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{1}],\ |\mathcal{N}_{i_{1}}[t^{\prime}+t_{1}]\cap\mathcal{L}_{\mathcal{M}}|\geq F+1\}$. By definitions of strong $(2F+1)$-robustness and $F$-local model, for all $t\geq t_{C}$ such that $|\mathcal{F}_{\mathcal{R}}[t]|>0$, there must exist $i_{1}\in\mathcal{F}_{\mathcal{R}}[t]\cap\mathcal{F}_{\mathcal{L}}$. Since $\mathcal{F}_{\mathcal{C}}$ is finite and $\mathcal{F}_{\mathcal{L}}\subseteq\mathcal{F}_{\mathcal{C}}$, there must exist a finite $T_{1}\geq 1$ such that $\mathcal{F}_{\mathcal{L}}\subseteq\bigcup_{t_{1}=0}^{T_{1}-1}\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{1}]$. Therefore, for any $t_{1}\in[0,T_{1}-1]$ where $|\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{1}]|>0$, $\mathcal{S}_{1,t_{1}}=\{i_{1}\in\mathcal{F}_{\mathcal{C}}\mid\mathcal{N}_{i_{1}}[t^{\prime}+t_{1}]\cap\mathcal{L}_{\mathcal{M}}|\geq F+1\}$ is nonempty. Then, through the BP-MSR algorithm, follower $i_{1}$ will use at least one $C_{r}$ to update $x_{i_{1}}[t^{\prime}+t_{1}+1]$.
@@
Furthermore, by definition, $\mathcal{G}_{\mathcal{R}}[t]$ is strongly $(2F+1)$-robust with respect to $\mathcal{L}\cup\mathcal{A}$ by Lemma 2 at $t$ if $|\mathcal{F}_{\mathcal{R}}[t]|>0$. Therefore, for any $t_{2}\in[T_{1},T_{2}-1]$ where $|\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{2}]|>0$, $\mathcal{S}_{2,t_{2}}=\{i_{2}\in\mathcal{S}_{X}(t^{\prime}+t_{2},\alpha^{t_{2}}\underline{\epsilon},\alpha^{t_{2}}\overline{\epsilon})\mid|\mathcal{D}_{i_{2}}[t^{\prime}+t_{2}]|\geq 2F+1\}$ is nonempty if $\mathcal{S}_{X}(t^{\prime}+t_{2},\alpha^{t_{2}}\underline{\epsilon},\alpha^{t_{2}}\overline{\epsilon})\neq\emptyset$. If it is nonempty, since $\mathcal{A}$ is $F$-local, $\mathcal{D}_{i_{2}}[t^{\prime}+t_{2}]$ contains at least $F+1$ normal agents $\forall i_{2}\in\mathcal{S}_{2,t_{2}}$. Then, by definition of $\mathcal{S}_{X}(t^{\prime}+t_{2},\alpha^{t_{2}}\underline{\epsilon},\alpha^{t_{2}}\overline{\epsilon})$, $x_{i_{2}}[t^{\prime}+t_{2}]>x_{j}^{i_{2}}[t^{\prime}+t_{2}]$ or $x_{i_{2}}[t^{\prime}+t_{2}]<x_{j}^{i_{2}}[t^{\prime}+t_{2}]$ $\forall j\in\mathcal{D}_{i_{2}}[t^{\prime}+t_{2}]$. This means $i_{2}$ will use at least one in-neighbor’s state within $\overline{m}[t^{\prime}]+\alpha^{t_{2}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{t_{2}}\overline{\epsilon}]$ to update $x_{i_{2}}[t^{\prime}+t_{2}+1]$, whose bounds are $x_{i_{2}}[t^{\prime}+t_{2}+1]\in[\overline{m}[t^{\prime}]+\alpha^{t_{2}+1}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{t_{2}+1}\overline{\epsilon}]$ by (8) and (9), which is also contained by $\overline{m}[t^{\prime}]+\alpha^{T_{2}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{T_{2}}\overline{\epsilon}]$. Now denote $\mathcal{S}_{2}=\cup_{t_{2}=T_{1}}^{T_{2}-1}\mathcal{S}_{2,t_{2}}$. Because (i) $\mathcal{S}_{2}\subseteq\mathcal{K}_{2}$, (ii) $\mathcal{K}_{1}\subseteq\mathcal{K}_{2}$, (iii) $\mathcal{S}_{2}\not\subseteq\mathcal{K}_{1}$, and (iv) $i_{2}\notin\mathcal{S}_{X}(t^{\prime}+T_{2},\alpha^{T_{2}}\underline{\epsilon},\alpha^{T_{2}}\overline{\epsilon})\ \forall i_{2}\in\mathcal{K}_{2}$, we get $|\mathcal{S}_{X}(t^{\prime}+T_{2},\alpha^{T_{2}}\underline{\epsilon},\alpha^{T_{2}}\overline{\epsilon})|<|\mathcal{S}_{X}(t^{\prime}+T_{1},\alpha^{T_{1}}\underline{\epsilon},\alpha^{T_{1}}\overline{\epsilon})|$.
@@
This logic of the existence of finite time $T_{p}\geq T_{p-1}+1$ such that $|\mathcal{S}_{X}(t^{\prime}+T_{p},\alpha^{T_{p}}\underline{\epsilon},\alpha^{T_{p}}\overline{\epsilon})|<|\mathcal{S}_{X}(t^{\prime}+T_{p-1},\alpha^{T_{p-1}}\underline{\epsilon},\alpha^{T_{p-1}}\overline{\epsilon})|$ can be continued iteratively for $p\geq 2$. We first define $\mathcal{K}_{p}=\{i_{p}\in\mathcal{F}_{\mathcal{R}}[t^{\prime}+T_{p}]\mid x_{i_{p}}[t^{\prime}+T_{p}]\in[\overline{m}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{T_{p}}\overline{\epsilon}]\}$. Furthermore, we define $\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]=\mathcal{N}_{i_{p}}[t^{\prime}+t_{p}]\cap\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{p}]\setminus\mathcal{S}_{X}(t^{\prime}+t_{p},\alpha^{t_{p}}\underline{\epsilon},\alpha^{t_{p}}\overline{\epsilon})$ for any $t_{p}\in[T_{p-1},T_{p}-1]$. Let $T_{p}\geq T_{p-1}+1$ be the earliest time where $\mathcal{K}_{p-1}\subseteq\mathcal{K}_{p}$. Using the prior arguments, we conclude that $x_{i_{p-1}}[t^{\prime}+T_{p}]\in[\overline{m}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{T_{p}}\overline{\epsilon}]$ $\forall i_{p-1}\in\mathcal{K}_{p-1}$, and thus $\mathcal{K}_{p-1}\subseteq\mathcal{K}_{p}$. Because all $i_{p-1}\in\mathcal{K}_{p-1}\subseteq\mathcal{F}_{\mathcal{C}}$ must belong to $\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{p}]$ at some $t_{p}\in[T_{p-1},T_{p}-1]$, such finite $T_{p}$ exists.
@@
Furthermore, using the prior arguments, for any $t_{p}\in[T_{p-1},T_{p}-1]$ where $|\mathcal{F}_{\mathcal{R}}[t^{\prime}+t_{p}]|>0$, we can have nonempty $\mathcal{S}_{p,t_{p}}=\{i_{p}\in\mathcal{S}_{X}(t^{\prime}+t_{p},\alpha^{t_{p}}\underline{\epsilon},\alpha^{t_{p}}\overline{\epsilon})\mid|\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]|\geq 2F+1\}$. If it is nonempty, since $\mathcal{A}$ is $F$-local, $\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]$ contains at least $F+1$ normal agents $\forall i_{p}\in\mathcal{S}_{p,t_{p}}$. Then, by definition of $\mathcal{S}_{X}(t^{\prime}+t_{p},\alpha^{t_{p}}\underline{\epsilon},\alpha^{t_{p}}\overline{\epsilon})$, $x_{i_{p}}[t^{\prime}+t_{p}]>x_{j}^{i_{p}}[t^{\prime}+t_{p}]$ or $x_{i_{p}}[t^{\prime}+t_{p}]<x_{j}^{i_{p}}[t^{\prime}+t_{p}]$ $\forall j\in\mathcal{D}_{i_{p}}[t^{\prime}+t_{p}]$. This means $i_{p}$ will use at least one in-neighbor’s state within $\overline{m}[t^{\prime}]+\alpha^{t_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{t_{p}}\overline{\epsilon}]$ to update $x_{i_{p}}[t^{\prime}+t_{p}+1]$, whose bounds are $x_{i_{p}}[t^{\prime}+t_{p}+1]\in[\overline{m}[t^{\prime}]+\alpha^{t_{p}+1}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{t_{p}+1}\overline{\epsilon}]$ with the similar reasoning as (8) and (9), which is also contained by $\overline{m}[t^{\prime}]+\alpha^{T_{p}}\underline{\epsilon},\overline{M}[t^{\prime}]-\alpha^{T_{p}}\overline{\epsilon}]$. Now denote $\mathcal{S}_{p}=\cup_{t_{p}=T_{p-1}}^{T_{p}-1}\mathcal{S}_{p,t_{p}}$. Then, because (i) $\mathcal{S}_{p}\subseteq\mathcal{K}_{p}$, (ii) $\mathcal{K}_{p-1}\subseteq\mathcal{K}_{p}$, (iii) $\mathcal{S}_{p}\not\subseteq\mathcal{K}_{p-1}$, and (iv) $i_{p}\notin\mathcal{S}_{X}(t^{\prime}+T_{p},\alpha^{T_{p}}\underline{\epsilon},\alpha^{T_{p}}\overline{\epsilon})\ \forall i_{p}\in\mathcal{K}_{p}$, we get $|\mathcal{S}_{X}(t^{\prime}+T_{p},\alpha^{T_{p}}\underline{\epsilon},\alpha^{T_{p}}\overline{\epsilon})|<|\mathcal{S}_{X}(t^{\prime}+T_{p-1},\alpha^{T_{p-1}}\underline{\epsilon},\alpha^{T_{p-1}}\overline{\epsilon})|$.
@@
$\overline{m}[t^{\prime}]+\alpha^{\overline{T}}\underline{\epsilon}\leq x_{i}[t^{\prime}+\overline{T}]\leq\overline{M}[t^{\prime}]-\alpha^{\overline{T}}\overline{\epsilon}.$
Considering $V[t]=\overline{M}[t]-\overline{m}[t]$, we get:
$V[t^{\prime}+\overline{T}]$ $=\overline{M}[t^{\prime}+\overline{T}]-\overline{m}[t^{\prime}+\overline{T}]$
$\leq\overline{M}[t^{\prime}]-\alpha^{\overline{T}}\overline{\epsilon}-\left(\overline{m}[t^{\prime}]+\alpha^{\overline{T}}\underline{\epsilon}\right)$
@@
Using the fact that $\underline{\epsilon}=C_{r}-\overline{m}[t^{\prime}]$ and $\overline{\epsilon}=\overline{M}[t^{\prime}]-C_{r}$, from (10), we get
@@
The proof of Theorem 1 follows the similar approach of [23, Theorem 1], but with a key distinction. While [23, Theorem 1] assumes that the entire graph is strongly $(2F+1)$-robust with respect to $\mathcal{L}$ over a fixed time period, our analysis assumes that some individual followers become active and thus belong to a strongly $(2F+1)$-robust subgraph infinitely often.
@@
The corollary shows that if there exists an infinite sequence of finite time instants $\{\tau_{i}\}_{i\in\mathbb{Z}_{\geq 0}}$ such that $\bigcup_{t=\tau_{i}}^{\tau_{i+1}}\mathcal{G}[t]$ has all normal followers activated via (1) (which implies that the union of graph is strongly $(2F+1)$-robust), then the normal followers can achieve full resilient leader-follower consensus. This result in a way generalizes [23, Theorem 1], which requires the entire network to be strongly $(2F+1)$-robust over every fixed window of $T$ time steps.